If m-9=9 and mn=36, find: m+n and m^2-n^2
Answer by rothauserc(4718)

m - 9 = 9
m = 18
m*n = 36
n = 36/18 = 2
(1) m + n = 18 + 2 = 20
(2) m^2 - n^2 = 18^2 - 2^2 = 320